# Enhancing Person-to-Person Virtual Try-On with Multi-Garment Virtual Try-Off

### v1

@@ 4 Experiments
Since VTOFF task is newly introduced, few open-source methods exist. We evaluate our proposed TryOffDiff model against the publicly available TryOffAnyone [37] to establish a baseline. To ensure reproducibility, we provide a detailed description of our experimental setup. We adopt DISTS [11] as the primary evaluation metric, supplemented by standard metrics for a comprehensive comparison. Qualitative results are also presented to showcase TryOffDiff’s ability to handle diverse and challenging inputs.
@@ 4.1 Experimental Setup
For person-to-person try-on, where ground-truth images are unavailable, we randomly pair garments across individuals and provide a text file of input image filenames for reproducibility. For Dress Code, a label (upper-body, lower-body, or dresses) indicates the garment to transfer; this is not needed for VITON-HD, which includes only upper-body garments.
@@
Denoising U-Net weights are initialized from Stable Diffusion v1.4 [28] and subsequently fine-tuned. The Adapter module and class Embedding layer are trained from scratch. The SigLIP encoder, VAE encoder, and VAE decoder are kept frozen throughout the training, which is performed end-to-end. Input reference images are padded along the width to achieve a square aspect ratio and resized to $512\times 512$ to match the pretrained SigLIP and VAE encoder’s format. Garment images undergo the same preprocessing during training. We use SigLIP-B/16-512 as image feature extractor, yielding 1,024 token embeddings of dimension 768, which the adapter–comprising a linear layer and normalization–reduces to $n=77$ conditioning embeddings of dimension $m=768$. For garment conditioning, we embed class labels (e.g. ‘upper-body’, ‘lower-body’, ‘dress’) using a trainable embedding layer. This layer has 3 classes and projects the labels into a 1,280-dimensional space, matching the dimension of the timestep embeddings. These class embeddings are then combined with the timestep embeddings via element-wise addition before being fed into the U-Net’s residual blocks.
@@
4.2 Ablation Studies
We perform ablation studies to assess key components of TryOffDiff, including the image encoder, adapter module, conditioning feature dimensions, and the impact of finetuning a pretrained U-Net versus training from scratch. The configurations for these experiments are summarized in Tab. 1. In addition, Fig. 4 illustrates the impact of guidance scale and inference steps on FID and DISTS scores, evaluated on the Dress Code-test set using the Euler noise scheduler. This experiment reveals how these hyperparameters influence reconstruction quality, providing insights into optimal inference setting for TryOffDiff.
@@ 4.2 Ablation Studies
Figure 4: Impact of guidance scale ($s$) and inference steps ($n$) on DISTS and FID scores. Experiments are conducted on DressCode-test with TryOffDiff using the Euler scheduler [20].
4.3 Quantitative Results
We evaluate TryOffDiff’s performance on the VITON-HD and Dress Code datasets, comparing it to baseline TryOffAnyone [37] and category-specific ablations. Results are presented in Tab. 1 (VITON-HD), Tab. 2 (Dress Code), and Tab. 3 (p2p-VTON).
[…]
Dresses
TryOffDiff-single 81.6 76.1 55.6 26.1 18.4 8.2 4.4 20.8
TryOffDiff-multi 81.6 76.1 55.5 26.4 18.9 8.0 4.6 21.2
Upper-body
TryOffDiff-single 80.8 73.8 47.8 31.6 17.1 5.2 4.7 21.6
TryOffDiff-multi 80.2 73.2 48.1 32.3 17.8 5.3 5.5 22.2
Lower-body
TryOffDiff-single 81.1 76.8 55.5 30.0 22.6 9.4 5.9 21.1
TryOffDiff-multi 80.4 75.6 54.1 30.8 22.4 9.0 6.8 21.7
All
TryOffDiff-merged 81.2 75.6 52.9 29.2 11.4 6.7 3.1 21.2
TryOffDiff-multi 78.6 75.4 57.4 25.0 12.8 6.8 4.1 21.3
Table 2: Quantitative comparison of multi-garment VTOFF on DressCode-test. We compare category-specific models (TryOffDiff-single) trained independently on dresses, upper-body, and lower-body garments against a class-conditioned model (TryOffDiff-multi) trained jointly across all garment types. Additionally, we report results for TryOffDiff-merged, which aggregates the outputs of the three category-specific models into a single set. While non-distributional metrics (e.g., SSIM, LPIPS) for merged results reflect simple averages across categories, distributional metrics (e.g., FID, KID) are computed over the merged output set, capturing holistic distributional properties.
VITON-HD Dress Code
[…]
CatVTON 12.0 3.5 3.9 8.4 1.9 3.1
OOTD + GT 10.8 2.8 2.0 7.5 3.4 2.5
OOTD + TryOffDiff 11.9 3.3 2.6 7.9 3.4 2.7
Table 3: Quantitative results for p2p-VTON. We evaluate OOTDiffusion (OOTD) with ground truth (GT) garments and TryOffDiff-predicted garments, alongside CatVTON, a specialized p2p-VTON model. Our TryOffDiff model, when integrated with a VTON model, achieves competitive performance, even though it is not explicitly trained for the p2p task.
The results in Tab. 1 reveal key insights into our design choices. First, finetuning a pretrained U-Net consistently outperforms training from scratch across all metrics, underscoring the advantage of leveraging pretrained weights for initializing the diffusion model. Second, the SigLIP encoder outperforms CLIP, especially in perceptual quality metrics like FID and DISTS, suggesting that SigLIP’s visual representations are more effective for detailed garment reconstruction. Finally, incorporating the adapter module further improves performance by more effectively conditioning the diffusion process with image features.
On VITON-HD dataset (Tab. 1), our method outperforms the baseline across all metrics except LPIPS, where TryOffAnyone achieves a better score, though TryOffDiff excels in global metrics like FID and DISTS which is our primary metric.
In our evaluation on the Dress Code dataset (Tab. 2), we consider: (i) category-specific TryOffDiff models trained individually on dresses, upper-body, and lower-body garments; (ii) a class-conditioned TryOffDiff model trained jointly on all categories using garment-type embeddings; and (iii) an aggregated result, denoted TryOffDiff-merged, where outputs from the three category-specific models are combined post hoc. The class-conditioned model remains competitive across all categories, demonstrating its robustness as a unified solution.
Figure 5: Qualitative comparison of garment-specific and multi-garment TryOffDiff. First row displays the conditioning image. Second row shows dataset-specific reconstructions produced by TryOffDiff model trained on individual datasets: VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12). Last row shows predictions of multi-garment TryOffDiff trained on full DressCode with garment type conditioning.
For person-to-person virtual try-on (Tab. 3), we integrate TryOffDiff with OOTDiffusion and compare it to CatVTON, a specialized p2p-VTON model. On VITON-HD, OOTD+TryOffDiff performs slightly better than CatVTON. On Dress Code, TryOffDiff, when paired with OOTD, delivers competitive performance, approaching specialized p2p-VTON models like CatVTON, especially in perceptual quality metrics.
These findings highlight TryOffDiff’s versatility and effectiveness, enabling high-quality garment reconstruction for both VTOFF and p2p-VTON tasks. Its improvements in structural similarity and perceptual quality translate into practical benefits, such as more realistic garment rendering in both VTOFF and p2p-VTON applications.
@@ 4.4 Qualitative Analysis
Figure 5 illustrates the performance of TryOffDiff in garment reconstruction, comparing two variants: garment-specific models (row 2) and multi-garment model (row 3). The conditioning images, displayed in first row, are sourced from VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12). Garment-specific models excel in precision, particularly for text and fine details (cols 2 and 5), likely due to the pretrained model initialization. In contrast, the multi-garment model offers versatility, delivering strong performance across diverse garment types and datasets, albeit with slight trade-offs in detail compared to specialized models. Nevertheless, it reconstructs a high level of detail (column 6) and shows robust generalization to VITON-HD, which was unseen during training. In some cases (cols 2, 7 and 9), it even produces better fidelity of color.
Figure 6 illustrates the results of p2p-VTON. Overall, neither method stands out as a definitive winner. CatVTON excels in preserving texture and pattern details but occasionally introduces diffusion artifacts (row 2). It can also transfer unintended attributes, such as skin color, from target model to the source model (row 1)–a limitation not observed in classical try-on models. OOTDiffusion combined with TryOffDiff avoids such attribute transfer issues but may struggle due to masking, resulting in residual source garment attributes in the output (rows 3 and 4).
These results highlight the potential of Virtual Try-Off task and the TryOffDiff model. Although TryOffDiff was not specifically trained for p2p-VTON, its integration with VTON models presents a promising approach, already demonstrating competitive performance compared to state-of-the-art person-to-person virtual try-on methods.
Figure 6: Qualitative comparison for Person-to-Person Virtual Try-On. Columns show: (a) person to be dressed, (b) CatVTON output using a person image with target garment as condition for direct p2p-VTON, (c) OOTDiffusion output with ground truth target garment, and (d) OOTDiffusion output with multi-garment TryOffDiff prediction.
@@ 5 Conclusion
In this paper we propose TryOffDiff, a latent diffusion model tailored to Virtual Try-Off. We substitute text conditioning mechanism with adapted SigLIP features to steer the generative process, enabling effective garment reconstructions. TryOffDiff achieves state-of-the art results on VITON-HD. Further experiments were conducted on DressCode, which contains three types of garments: upper-body, lower-body and dresses. While the baseline architecture can be trained separately for each category, we enhance TryOffDiff by introducing class conditioning. This advancement allows a single model to perform multi-garment reconstruction, successfully capturing all garment types with minimal loss in detail quality. To our knowledge, this is the first method to address multi-garment VTOFF in a unified framework.
Moreover, we demonstrate that integrating VTOFF with VTON models enables Person-to-Person Virtual Try-On, achieving performance comparable to specialized models. This pipeline offers practical advantages: it leverages existing VTON models for a wider applications and decouples garment reconstruction from other generative artifacts, simplifying the assessment of p2p-VTON outcomes.
Despite these promising results, challenges persist in preserving complex structures, such as logos and printed designs. Future work could explore advanced generative models, alternative visual conditioning methods and additional losses to enhance detail preservation.

### v2

@@ 4 Experiments
Since Virtual Try-Off is a newly introduced task, few open‐source methods exist. We evaluate our Multi-Garment TryOffDiff (MGT) against two publicly-available baselines: TryOffDiff [32] and TryOffAnyone [35], to establish benchmarks. We adopt DISTS [11] as our primary metric and report additional standard measures for a comprehensive evaluation. We also present qualitative examples to illustrate MGT’s performance on diverse inputs.
@@ 4.1 Experimental Setup
For person-to-person (p2p) try-on, where ground-truth are unavailable, we randomly pair garments across individuals and provide a text file of input image filenames for reproducibility. For DressCode, a label (upper-body, lower-body, or dresses) indicates the garment to transfer; this is not needed for VITON-HD, which includes only upper-body garments.
@@
Denoising U-Net weights are initialized from Stable Diffusion v1.4 [27] and subsequently finetuned. The Adapter module and class Embedding layer are trained from scratch. The SigLIP encoder, VAE encoder, and VAE decoder are kept frozen throughout the training, which is performed end-to-end. Input reference images are padded along the width to achieve a square aspect ratio and resized to $512\times 512$ to match the pretrained SigLIP and VAE encoder’s format. Garment images undergo the same preprocessing during training. We use SigLIP-B/16-512 as image feature extractor, yielding 1,024 token embeddings of dimension 768, which the adapter–comprising a linear layer and normalization–reduces to $n=77$ conditioning embeddings of dimension $m=768$. For garment conditioning, we embed class labels (e.g. ‘upper-body’, ‘lower-body’, ‘dress’) using a trainable embedding layer. This layer has 3 classes and projects the labels into a 1,280-dimensional space, matching the dimension of the timestep embeddings. These class embeddings are then combined with the timestep embeddings via element-wise addition before being fed into the U-Net’s residual blocks.
@@
4.2 Quantitative Results
We compare MGT to TryOffDiff [32] and TryOffAnyone [35] on the VITON-HD and DressCode datasets, with results reported in Table 1 (DressCode), Table 2 (VITON-HD), and Table 3 (p2p-VTON). TryOffDiff and TryOffAnyone are trained on VITON-HD (upper-body only), while MGT is trained on DressCode, supporting multiple garment categories (upper-body, lower-body, dresses).
[…]
Upper-body
TryOffDiff-single 80.8 73.8 47.8 31.6 17.1 5.2 4.7 21.6
MGT (ours) 80.2 73.2 48.1 32.3 17.8 5.3 5.5 22.2
Lower-body
TryOffDiff-single 81.1 76.8 55.5 30.0 22.6 9.4 5.9 21.1
MGT (ours) 80.4 75.6 54.1 30.8 22.4 9.0 6.8 21.7
Dresses
TryOffDiff-single 81.6 76.1 55.6 26.1 18.4 8.2 4.4 20.8
MGT (ours) 81.6 76.1 55.5 26.4 18.9 8.0 4.6 21.2
Table 1: Quantitative results on DressCode-test. Comparing TryOffDiff(single-category) to MGT(unified model).
No prior VTOFF model handles multiple garment categories. Therefore, to establish a baseline and to measure the cost of unified modeling, we train three category-specific TryOffDiff variants on DressCode (upper-body, lower-body, dresses) and compare them against MGT, a single model conditioned on garment type. This setup tests whether supporting multiple categories in a unified architecture results in performance degradation. Table 1 shows MGT achieves comparable performance across all categories, with minor trade-offs (e.g., DISTS: 22.2 vs. 21.6 for upper-body, LPIPS: 30.8 vs. 30.0 for lower-body, and FID: 18.9 vs. 18.4 for dresses). This shows that a single, unified approach can replace multiple specialized models without significant performance loss.
[…]
TryOffDiff [32] 79.5 70.4 46.2 32.4 25.1 9.4 8.9 23.0
TryOffAnyone [35] 71.9 - - 17.2 25.3 - 2.0 21.0
[…]
We further assess MGT’s generalization in a cross-dataset setting by evaluating it on VITON-HD, despite being trained exclusively on DressCode. As shown in Table 2, MGT outperforms TryOffDiff in FID (21.9 vs. 25.1) and remains competitive across other metrics. This indicates strong generalization despite the domain shift.
VITON-HD DressCode
[…]
CatVTON [10] 12.0 3.5 3.9 8.4 1.9 3.1
OOTD [39] + Ground truth 10.8 2.8 2.0 7.5 3.4 2.5
OOTD [39] + MGT (ours) 11.9 3.3 2.6 7.9 3.4 2.7
Table 3: Quantitative results for p2p-VTON. We evaluate OOTDiffusion (OOTD) with ground truth (GT) garments and MGT-predicted garments, alongside CatVTON, a specialized p2p-VTON model. Our model’s output, when integrated with a VTON model, achieves competitive performance, even though it is not explicitly trained for the p2p task.
In p2p-VTON setup (Tab. 3), we pair MGT with OOTDiffusion and compare the results to CatVTON, a dedicated p2p-VTON model. On VITON-HD, OOTD+MGT slightly outperforms CatVTON. On DressCode, MGT achieves competitive results, especially in perceptual metrics, showing its effectiveness in tasks beyond its training objective.
These findings highlight MGT’s versatility across both VTOFF and p2p-VTON tasks. Its improvements in perceptual quality and structural fidelity translate to practical benefits in realistic garment rendering.
Lastly, Figure 4 explores the influence of inference hyperparameters (guidance scale and number of steps) on FID and DISTS scores. This analysis provides insights into optimal inference configurations for MGT.
@@ 4.2 Quantitative Results
Figure 4: Impact of guidance scale ($s$) and inference steps ($n$) on DISTS and FID scores. Evaluated on DressCode-test with MGT using the Euler scheduler [20].
Figure 5: Qualitative comparison of garment-specific and Multi-Garment TryOffDiff. First row displays the reference image. Second row shows dataset-specific reconstructions produced by TryOffDiff model trained on single category: VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12). Last row shows predictions of Multi-Garment TryOffDiff trained on full DressCode with garment type conditioning.
(a) Input
(b) ToD [32]
(c) ToA [35]
(d) MGT
(e) GT
Figure 6: Qualitative comparison on VITON-HD. Despite not being trained on VITON-HD, MGT achieves competitive results compared to baselines specifically trained on it: TryOffDiff (ToD) [32] and TryOffAnyone (ToA) [35].
@@ 4.3 Qualitative Analysis
Figure 5 shows outputs from two sets of models: category-specific TryOffDiff variants (row 2) and our unified MGT model (row 3). Top row shows input images drawn from; VITON-HD (cols 1–3), DressCode upper-body (4–6), lower-body (7–9), and dresses (10–12). While all models produce visually plausible outputs, MGT matches or surpasses single-category models in several cases. It maintains clear text (cols 1, 6), emblems (col 4), and accurate garment shapes (col 5), and recovers detailed textures (cols 2, 10). MGT also performs strongly on the unseen VITON-HD dataset (cols 1–3), demonstrating robust generalization across garment types and domains.
To complement the quantitative results on the VITON-HD (Tab. 2), we provide a qualitative comparison in Figure 6. Despite not being trained on VITON-HD, MGT produces outputs that are visually competitive with baselines that were trained specifically on this dataset, including TryOffDiff [32] and TryOffAnyone [35]. It maintains structural coherence and texture detail, even under challenging conditions involving complex patterns or non-frontal poses. These findings affirm MGT’s ability to generalize without the need for per-category or per-dataset tuning.
Figure 7 shows results for p2p-VTON. CatVTON transfers textures and patterns accurately but occasionally introduces artifacts (rows 1–2) or unintended changes such as skin tone shifts (row 1). In contrast, combining OOTDiffusion with MGT yields consistent, artifact-free outputs. Although no method dominates in all scenarios, the decoupling of garment reconstruction (via MGT) from rendering (via VTON) results in more interpretable and stable outputs.
(a) Single Garment
(b) Multi-Garment
Figure 7: Qualitative comparison for Person-to-Person Virtual Try-On in single and multi-garment settings. Columns: (i) source model image (person to be dressed), (ii) output from CatVTON, using a person image with target garments as condition for direct p2p-VTON, (iii) output from OOTDiffusion (OOTD), conditioned on the ground-truth (GT) target garments, and (iv) output from OOTDiffusion with our Multi-Garment TryOffDiff (MGT) pipeline. Each result shows the generated outfit on the source model, with the corresponding target garments shown in the top-left and top-right corners. For single-garment settings (a), only the top-left garment is applied. For multi-garment settings (b), garments are applied sequentially: the top-left garment is applied first, and its output serves as input for applying the top-right garment.
@@ 5 Conclusion
We introduced Multi-Garment TryOffDiff (MGT), a unified latent diffusion model for the Virtual Try-Off (VTOFF) task. By incorporating a lightweight class-conditioning mechanism, MGT supports the reconstruction of diverse garment types—upper-body, lower-body, and dresses—from a single reference image. Unlike prior work, which relies on separate models for each garment category, MGT achieves near-specialized performance within a single framework. On the DressCode dataset, it matches or surpasses category-specific TryOffDiff variants while supporting multiple categories simultaneously. To our knowledge, MGT is the first unified model to address multi-garment VTOFF. Moreover, MGT demonstrates strong cross-dataset generalization. Despite being trained exclusively on DressCode, it achieves competitive results on VITON-HD, highlighting its robustness to domain shifts and unseen garments. Beyond VTOFF, we show that MGT can be seamlessly integrated into person-to-person Virtual Try-On (p2p-VTON) pipelines. When paired with OOTDiffusion, it achieves performance on par with specialized p2p models like CatVTON, while offering a modular and interpretable pipeline, separating garment reconstruction from person rendering.
Limitations and Future Work. MGT currently supports only three garment categories and does not handle layered clothing, constrained by the available dataset annotations. Fine-grained texture recovery and logo preservation also remain challenging, particularly for complex garments. Future research could focus on expanding garment coverage, refining conditioning mechanisms, and exploring higher-capacity architectures or perceptual training objectives to improve detail fidelity.
In summary, MGT marks a step toward general-purpose virtual try-off by unifying garment reconstruction across categories in a single, adaptable model. Its strong performance, generalization capabilities, and modularity make it a valuable component for future virtual fashion systems.